# Seepage analysis using a polygonal cell-based smoothed finite element method

### v1

@@ Abstract
This work develops a polygonal cell-based smoothed finite element method for steady-state, transient, and free-surface seepage in saturated porous media. Wachspress interpolation on convex polygonal elements is combined with cell-based gradient smoothing, so that element matrices are assembled using boundary integrals without in-element derivatives. Polygonal, quadtree, and hybrid quadtree–polygonal meshes are employed to accommodate local refinement and hanging nodes, and a solution-driven adaptive strategy further concentrates resolution near steep gradients and wet–dry transitions. Free-surface seepage is solved using a fixed-mesh iterative scheme that updates the wetted region, permeability field, and boundary conditions. Benchmark tests demonstrate accurate hydraulic-head and free-surface predictions, and show that adaptivity attains similar accuracy with substantially fewer degrees of freedom and CPU time.
@@
Seepage analysis , Smoothed finite element method , Polygonal elements , Wachspress basis functions , Hybrid quadtree meshes , Free-surface flow
@@
[inst3]organization=Yunnan Institute of Water Resources and Hydropower Research,city=Kunming, postcode=650228, state=Yunnan, country=China
{highlights}
A polygonal cell-based smoothed finite element method is developed for seepage analysis in porous media.
Wachspress interpolation on polygonal elements is combined with gradient smoothing requiring boundary integrals only.
Polygonal, quadtree, and hybrid discretizations naturally accommodate hanging nodes and local refinements without constraints.
A fixed-mesh iterative procedure is implemented to efficiently capture free surfaces in seepage problems.
Adaptive refinement achieves comparable accuracy with markedly reduced degrees of freedom and CPU time.
@@ 1 Introduction
Despite its broad applicability, conventional FEM formulations based on triangular or quadrilateral elements face several limitations. Triangular meshes are straightforward to generate but often exhibit lower accuracy, while quadrilateral meshes require more manual control [33, 9]. Standard isoparametric elements are also sensitive to mesh distortion, since the Jacobian must remain positive and element angles must satisfy geometric constraints [25, 46]. Overall, these drawbacks make the conventional FEM less robust and less efficient for problems with complex geometries and highly distorted meshes.
To alleviate the above mesh-sensitivity issues while retaining the simplicity of low-order discretizations, the smoothed finite element method (SFEM) introduces a strain/gradient smoothing operation over appropriately constructed smoothing domains, in which gradients are averaged and—via the divergence theorem—evaluated through boundary integrals in the physical space. This treatment weakens the dependence on the inverse Jacobian inherent to isoparametric mappings, thereby relaxing geometric restrictions and improving numerical robustness on distorted meshes [23, 24, 41]. Existing studies have further shown that SFEM can achieve higher accuracy and faster energy-norm convergence than the conventional FEM [23, 37]. Consequently, SFEM is less sensitive to element distortion and can accommodate elements of arbitrary geometry, which is particularly appealing for seepage-flow analyses in complex geometries where locally steep hydraulic gradients and evolving wet–dry interfaces can exacerbate mesh-quality issues.
@@
While CSFEM improves robustness through gradient smoothing, the overall flexibility and mesh quality of the discretization remain crucial for seepage simulations in complex geometries, which has motivated the increasing use of polygonal and polyhedral discretizations to alleviate meshing constraints and enhance robustness on locally refined or highly irregular grids. For example, polygonal scaled boundary formulations have been implemented in commercial FE platforms and applied to steady-state and transient seepage analyses on polygon/quadtree meshes [39, 38]. Beyond scaled-boundary approaches, conservative variational schemes have been developed directly on general polygonal meshes for Darcy flow, including lowest-order weak Galerkin methods with optimal convergence and continuous normal fluxes [26], as well as mimetic finite difference discretizations designed for diffusion/Darcy operators on unstructured polygonal meshes and nonmatching local refinements [19]. More recently, virtual element discretizations have provided another flexible route for Darcy flow on polygonal meshes, with rigorous analysis and extensions to fractured porous media [35]. Overall, these studies indicate that polygonal elements can offer tangible advantages for seepage modelling; coupling such geometric flexibility with the CSFEM smoothing framework therefore provides a promising route to simultaneously enhance discretization versatility and numerical robustness in the present polygonal CSFEM approach.
@@
The objective of this work is to develop a polygonal CSFEM for seepage analysis in saturated porous media, using Wachspress interpolation on convex and nonconvex polygonal elements and gradient smoothing on triangular subcells so that smoothed hydraulic head gradients are obtained from boundary integrals only. Hybrid quadtree–polygonal meshes are employed to handle hanging nodes without additional constraint equations, and a fixed-mesh iterative procedure is adopted for free-surface seepage. The proposed approach is validated through benchmark problems for steady-state, transient, and free-surface seepage, showing close agreement with analytical and reference solutions and demonstrating strong robustness on both polygonal and quadtree meshes. The remainder of this paper is organized as follows: Section 2 presents the polygonal CSFEM formulation, Sections 3 and 4 describe the Wachspress interpolation and free-surface strategy, Section 5 reports numerical validations, and Section 6 concludes the paper.
@@ 2 Theoretical formulation of the polygonal CSFEM for seepage problems
We consider a saturated porous medium occupying the domain $\Omega$, bounded by $\Gamma=\Gamma_{h}\cup\Gamma_{q}$ with $\Gamma_{h}\cap\Gamma_{q}=\varnothing$, where $\Gamma_{h}$ and $\Gamma_{q}$ are the portions of the boundary on which the hydraulic head and the normal flux are prescribed, respectively. Fluid motion follows Darcy’s law together with the mass conservation principle. The transient seepage problem is governed by
@@
The boundary conditions consist of prescribed-head and prescribed-flux boundaries:
@@
where $\bar{h}$ is the prescribed hydraulic head on $\Gamma_{h}$, $\bar{q}$ is the prescribed outward normal flux on $\Gamma_{q}$, and $\mathbf{n}$ is the outward unit normal.
To arrive at the weak form, let $w$ be an admissible weighting function with $w=0$ on $\Gamma_{h}$. Multiplying Eq. (1) by $w$ and integrating over $\Omega$ gives
$\int_{\Omega}w\,\nabla\cdot(\mathbf{k}\nabla h)\,d\Omega+\int_{\Omega}w\,p\,d\Omega-\int_{\Omega}w\,S_{s}\,\dot{h}\,d\Omega=0.$ (4)
Using the divergence theorem for the first term:
@@
With $w=0$ on $\Gamma_{h}$ and the Neumann condition enforced on $\Gamma_{q}$, the Galerkin weak form of the governing equation can be written as
$\int_{\Omega}(\nabla w)^{T}\mathbf{k}\nabla h\,d\Omega+\int_{\Omega}w\,S_{s}\,\dot{h}\,d\Omega=\int_{\Omega}w\,p\,d\Omega+\int_{\Gamma_{q}}w\,\bar{q}\,d\Gamma.$ (6)
2.1 Smoothed gradient in a polygonal element
Let a polygonal element $\Omega_{e}$ be subdivided into $n_{c}$ non-overlapping smoothing cells $\{\Omega_{C}\}_{C=1}^{n_{c}}$ with boundaries $\Gamma_{C}$ and areas $A_{C}$. Within each $\Omega_{C}$, the smoothed gradient of $h$ is defined as
$\widetilde{\nabla}h=\frac{1}{A_{C}}\int_{\Omega_{C}}\nabla h(\mathbf{x})\,d\Omega.$ (7)
Applying the divergence theorem to the vector field $h\mathbf{e}_{i}$ yields
$\widetilde{\nabla}h=\frac{1}{A_{C}}\int_{\Gamma_{C}}h(\mathbf{x})\,\mathbf{n}(\mathbf{x})\,d\Gamma,$ (8)
which removes the need to evaluate derivatives of shape functions inside $\Omega_{C}$.
Let $h(\mathbf{x})=\sum_{I=1}^{n}N_{I}(\mathbf{x})h_{I}$. Substituting into Eq. (8) gives
$\widetilde{\nabla}h=\sum_{I=1}^{n}\widetilde{\mathbf{B}}_{I}^{(C)}\,h_{I},$ (9)
@@ 2.1 Smoothed gradient in a polygonal element
If $\Gamma_{C}$ consists of straight edges $e=1,\dots,m_{C}$ with length $L_{e}$ and constant normal $\mathbf{n}_{e}$, then
@@
The edge integral may be evaluated using one-dimensional Gauss quadrature:
$\int_{e}N_{I}(\mathbf{x})\,d\Gamma\approx\sum_{g=1}^{n_{g}}N_{I}(\mathbf{x}_{g})\,w_{g}\,L_{e}.$ (12)
2.2 Element stiffness and capacity matrices
With $\widetilde{\nabla}h$ replacing the compatible gradient $\nabla h$ in the weak form, the smoothed element stiffness matrix becomes
$\mathbf{K}_{e}=\sum_{C=1}^{n_{c}}A_{C}\,\widetilde{\mathbf{B}}^{(C)T}\mathbf{k}\,\widetilde{\mathbf{B}}^{(C)},$ (13)
where
$\widetilde{\mathbf{B}}^{(C)}=\begin{bmatrix}\widetilde{\mathbf{B}}_{1}^{(C)}&\widetilde{\mathbf{B}}_{2}^{(C)}&\cdots&\widetilde{\mathbf{B}}_{n}^{(C)}\end{bmatrix}$. (14)
The capacity matrix follows the standard interpolation of $h$:
$\mathbf{M}_{e}=\int_{\Omega_{e}}N^{T}S_{s}N\,d\Omega=\sum_{C=1}^{n_{c}}\int_{\Omega_{C}}N^{T}S_{s}N\,d\Omega.$ (15)
@@ 2.2 Element stiffness and capacity matrices
Assembly over all elements leads to the global semi-discrete system
$\mathbf{K}\mathbf{H}+\mathbf{M}\dot{\mathbf{H}}=\mathbf{F},$ (18)
where $\mathbf{F}=\mathbf{f}_{p}+\mathbf{f}_{q}$.
For the fully implicit backward Euler scheme with time step $\Delta t$, the time derivative is approximated by
$\dot{\mathbf{H}}^{\,n+1}\approx\frac{\mathbf{H}^{n+1}-\mathbf{H}^{n}}{\Delta t}.$ (19)
Substituting into the semi-discrete system (18) and evaluating the operators at $t^{n+1}$ gives
@@ 3 Polygonal interpolation using Wachspress basis functions
Consider a convex $n$-gon with ordered vertices $\{\mathbf{x}_{k}\}_{k=1}^{n}$ in the physical plane. For each vertex $k$, the Wachspress weight function $w_{k}(\mathbf{x})$ is defined by
@@
where $f_{1}$ and $f_{2}$ denote the two polygon edges adjacent to vertex $k$, $\mathbf{n}_{f_{i}}$ are their outward unit normals, and
$h_{f_{i}}(\mathbf{x})=\big(\mathbf{x}_{k}-\mathbf{x}\big)\cdot\mathbf{n}_{f_{i}}$ (22)
are the corresponding signed distances from $\mathbf{x}$ to the edges.
The Wachspress shape function associated with vertex $k$ is then given by
@@
The Wachspress basis forms a nonnegative, $C^{0}$-continuous, linearly complete interpolation on convex polygons and satisfies the partition-of-unity property
@@
In the polygonal CSFEM, Wachspress functions enter the formulation only through boundary integrals of the type
@@
and their edge-based representation
@@
which involve only $N_{I}$ but not $\nabla N_{I}$. Because $N_{I}$ is evaluated along straight edges, the rational Wachspress functions can be integrated accurately by low-order Gauss quadrature.

### v2

# Two-dimensional seepage analysis using a polygonal cell-based smoothed finite element method

@@ Abstract
This study develops a polygonal cell-based smoothed finite element method (CSFEM) for two-dimensional seepage analyses in porous media, covering steady-state, transient, and free-surface problems. Wachspress interpolation on convex polygonal elements is combined with cell-based gradient smoothing, so that element matrices are assembled using boundary integrals only, avoiding in-element derivatives and improving robustness on distorted and locally refined meshes. To improve efficiency, a solution-driven adaptive refinement strategy is employed to concentrate resolution near steep hydraulic gradients and evolving wet–dry interfaces. Free-surface seepage is handled by a fixed-mesh iterative scheme that updates the wetted region and boundary conditions to track the phreatic surface. Benchmark tests validate the formulation against analytical solutions and high-fidelity FEM references. In steady seepage examples, the proposed polygonal CSFEM reproduces linear hydraulic-head fields to near machine precision and yields smaller head errors than conventional FEM at the same characteristic mesh size. In transient problems, accurate head evolution and stable time responses are obtained, while adaptive refinement efficiently resolves localized high-gradient zones. For free-surface cases, the method captures the phreatic-surface profile and seepage-face development reliably without remeshing. The quadtree refinement and adaptivity provide substantial efficiency gains in degrees of freedom and runtime for a prescribed accuracy level.
@@
Seepage analysis , Smoothed finite element method , Polygonal elements , Wachspress basis functions , Free-surface flow , Adaptive refinement
@@
[inst3]organization=College of Harbour, Coastal and Offshore Engineering,city=Nanjing, postcode=210098, state=Jiangsu, country=China
[…]
[inst4]organization=Yunnan Institute of Water Resources and Hydropower Research,city=Kunming, postcode=650228, state=Yunnan, country=China
@@ 1 Introduction
Despite its broad applicability, conventional FEM formulations based on triangular or quadrilateral elements face several limitations. Triangular meshes are straightforward to generate but may exhibit lower accuracy, whereas quadrilateral meshes often require more manual control [38, 9]. Standard isoparametric elements are also sensitive to mesh distortion, since the Jacobian must remain positive and element angles must satisfy geometric constraints [27, 52]. Overall, these drawbacks make the conventional FEM less robust and less efficient for problems with complex geometries and highly distorted meshes.
To alleviate the above mesh-sensitivity issues while retaining the simplicity of low-order discretizations, the smoothed finite element method (SFEM) introduces a gradient smoothing operation over appropriately constructed smoothing domains, in which gradients are averaged and—via the divergence theorem—evaluated through boundary integrals in the physical space. This treatment reduces the dependence on the inverse Jacobian inherent in isoparametric mappings, thereby relaxing geometric restrictions and improving robustness on distorted meshes [25, 26, 47]. Existing studies have further shown that SFEM can achieve higher accuracy and faster energy-norm convergence than the conventional FEM [25, 43]. Consequently, SFEM is less sensitive to element distortion and can accommodate elements of arbitrary geometry, which is particularly appealing for seepage-flow analyses in complex geometries where locally steep hydraulic gradients and evolving wet–dry interfaces can exacerbate mesh-quality issues.
@@
While CSFEM improves robustness through gradient smoothing, the overall flexibility and mesh quality of the discretization remain crucial for seepage simulations in complex geometries, which has motivated the increasing use of polygonal discretizations to alleviate meshing constraints and enhance robustness on locally refined or highly irregular grids. For example, polygonal scaled boundary formulations have been implemented in commercial FE platforms and applied to steady-state and transient seepage analyses on polygon/quadtree meshes [45, 44]. Beyond scaled-boundary approaches, conservative variational schemes have been developed directly on general polygonal meshes for Darcy flow, including lowest-order weak Galerkin methods with optimal convergence and continuous normal fluxes [29], as well as mimetic finite difference discretizations designed for diffusion/Darcy operators on unstructured polygonal meshes and nonmatching local refinements [21]. More recently, virtual element discretizations have provided another flexible route for Darcy flow on polygonal meshes, with rigorous analysis and extensions to fractured porous media [41]. Overall, these studies indicate that polygonal elements can offer tangible advantages for seepage modelling; coupling such geometric flexibility with the CSFEM smoothing framework therefore provides a promising route to simultaneously enhance discretization versatility and numerical robustness in the present polygonal CSFEM approach.
@@
The objective of this work is to develop a polygonal CSFEM for seepage analysis in saturated porous media, using Wachspress interpolation on convex polygonal elements and gradient smoothing on triangular subcells so that smoothed hydraulic head gradients are obtained from boundary integrals only. Hybrid quadtree meshes are employed to handle hanging nodes without additional constraint equations, and a fixed-mesh iterative procedure is adopted for free-surface seepage. The proposed approach is validated through benchmark problems for steady-state, transient, and free-surface seepage. The remainder of this paper is organized as follows: Section 2 presents the polygonal CSFEM formulation, Sections 3 and 4 describe the Wachspress interpolation and free-surface strategy, Section 5 reports numerical validations, and Section 6 concludes the work.
@@ 2 Theoretical formulation of the polygonal CSFEM for seepage problems
We consider a porous medium occupying the domain $\Omega\subset\mathbb{R}^{2}$, bounded by $\Gamma=\Gamma_{h}\cup\Gamma_{q}$ with $\Gamma_{h}\cap\Gamma_{q}=\varnothing$, where $\Gamma_{h}$ and $\Gamma_{q}$ are the portions of the boundary on which the hydraulic head and the outward normal flux are prescribed, respectively. The transient seepage problem is described by Darcy’s law and mass conservation, leading to the diffusion-type governing equation [18]
@@
The boundary conditions consist of prescribed-head and prescribed-flux boundaries
@@
where $\bar{h}$ is the prescribed hydraulic head on $\Gamma_{h}$, $\bar{q}$ is the prescribed outward normal flux on $\Gamma_{q}$, and $\mathbf{n}$ is the outward unit normal. For transient analyses, the initial condition is specified as
[…]
2.1 Galerkin weak form and spatial discretization
Let $w$ be an admissible weighting function such that $w=0$ on $\Gamma_{h}$. Multiplying Eq. (1) by $w$ and integrating over $\Omega$ gives
$\int_{\Omega}w\,\nabla\cdot(\mathbf{k}\nabla h)\,d\Omega+\int_{\Omega}w\,p\,d\Omega-\int_{\Omega}w\,S_{s}\,\frac{\partial h}{\partial t}\,d\Omega=0.$ (5)
Applying the divergence theorem to the first term yields
@@ 2.1 Galerkin weak form and spatial discretization
With $w=0$ on $\Gamma_{h}$ and using the Neumann condition on $\Gamma_{q}$, the Galerkin weak form can be written as
$\int_{\Omega}(\nabla w)^{T}\mathbf{k}\nabla h\,d\Omega+\int_{\Omega}w\,S_{s}\,\frac{\partial h}{\partial t}\,d\Omega=\int_{\Omega}w\,p\,d\Omega+\int_{\Gamma_{q}}w\,\bar{q}\,d\Gamma.$ (7)
Let $\Omega$ be discretized into polygonal elements $\{\Omega_{e}\}$. The trial and test functions are approximated by
$h^{h}(\mathbf{x},t)=\sum_{I=1}^{n_{e}}N_{I}(\mathbf{x})\,h_{I}(t),\qquad w^{h}(\mathbf{x})=\sum_{J=1}^{n_{e}}N_{J}(\mathbf{x})\,\delta h_{J},$ (8)
where $N_{I}$ are polygonal shape functions (e.g., Wachspress interpolants on convex polygons), $n_{e}$ is the number of nodes of $\Omega_{e}$, and $\delta h_{J}$ are arbitrary nodal variations satisfying essential boundary conditions. Substituting Eq. (8) into Eq. (7) and assembling over all elements leads to the standard semi-discrete system
$\mathbf{K}\mathbf{H}+\mathbf{M}\dot{\mathbf{H}}=\mathbf{F},$ (9)
where $\mathbf{H}$ collects the global nodal heads, $\dot{\mathbf{H}}$ is its time derivative, and $\mathbf{F}$ includes source and Neumann contributions.
2.2 Smoothed gradient over polygonal smoothing cells
For each polygonal element $\Omega_{e}$, a cell-based smoothing strategy is employed. The element is subdivided into $n_{c}$ non-overlapping smoothing cells $\{\Omega_{C}\}_{C=1}^{n_{c}}$ such that $\Omega_{e}=\cup_{C=1}^{n_{c}}\Omega_{C}$, with boundaries $\Gamma_{C}$ and areas $A_{C}$. A piecewise-constant smoothing function is defined by [28]
$\Phi(\mathbf{x}-\mathbf{x}_{C})=\begin{cases}1/A_{C},&\mathbf{x}\in\Omega_{C},\\ 0,&\mathbf{x}\notin\Omega_{C},\end{cases}$ (10)
where $\mathbf{x}_{C}$ is a representative point of the smoothing cell. The smoothed gradient of the hydraulic head is defined as the weighted-average of the compatible gradient
$\widetilde{\nabla}h(\mathbf{x}_{C})=\int_{\Omega_{C}}\nabla h(\mathbf{x})\,\Phi(\mathbf{x}-\mathbf{x}_{C})\,d\Omega.$ (11)
Applying integration by parts to Eq. (11) gives
$\widetilde{\nabla}h(\mathbf{x}_{C})=\int_{\Gamma_{C}}h(\mathbf{x})\,\mathbf{n}(\mathbf{x})\,\Phi(\mathbf{x}-\mathbf{x}_{C})\,d\Gamma-\int_{\Omega_{C}}h(\mathbf{x})\,\nabla\Phi(\mathbf{x}-\mathbf{x}_{C})\,d\Omega.$ (12)
Substituting Eq. (10) into Eq. (12), and noting that $\nabla\Phi(\mathbf{x}-\mathbf{x}_{C})=\mathbf{0}$ inside $\Omega_{C}$ for the constant smoothing function, we obtain the boundary-integral form
$\widetilde{\nabla}h(\mathbf{x}_{C})=\frac{1}{A_{C}}\int_{\Gamma_{C}}h(\mathbf{x})\,\mathbf{n}(\mathbf{x})\,d\Gamma,$ (13)
where $\mathbf{x}_{C}$ is a representative point of the smoothing cell, $\Gamma_{C}=\partial\Omega_{C}$ is its boundary, and $\mathbf{n}(\mathbf{x})$ is the outward unit normal on $\Gamma_{C}$. This boundary-integral representation involves only values of $h$ on $\Gamma_{C}$, so no in-cell evaluation of shape-function derivatives $\nabla N_{I}$ is required.
Using the interpolation $h^{h}(\mathbf{x})=\sum_{I=1}^{n_{e}}N_{I}(\mathbf{x})h_{I}$, Eq. (13) yields
$\widetilde{\nabla}h^{h}(\mathbf{x}_{C})=\sum_{I=1}^{n_{e}}\widetilde{\mathbf{B}}_{I}^{(C)}\,h_{I},$ (14)
@@ 2.2 Smoothed gradient over polygonal smoothing cells
For straight boundary segments $e=1,\dots,m_{C}$ of $\Gamma_{C}$ with length $L_{e}$ and constant outward normal $\mathbf{n}_{e}$, one obtains
@@
The edge integral can be evaluated by one-dimensional Gauss quadrature
$\int_{e}N_{I}(\mathbf{x})\,d\Gamma\approx\sum_{g=1}^{n_{g}}N_{I}(\mathbf{x}_{g})\,w_{g}\,L_{e},$ (17)
where $\{\mathbf{x}_{g},w_{g}\}$ are Gauss points and weights on the edge.
2.3 Element matrices with smoothed gradients
In the Galerkin setting, both the trial and test function gradients are replaced by their smoothed counterparts over each smoothing cell. Defining the cell-wise smoothed gradient matrix
$\widetilde{\mathbf{B}}^{(C)}=\begin{bmatrix}\widetilde{\mathbf{B}}_{1}^{(C)}&\widetilde{\mathbf{B}}_{2}^{(C)}&\cdots&\widetilde{\mathbf{B}}_{n_{e}}^{(C)}\end{bmatrix}$, (18)
the smoothed element stiffness matrix is obtained as
$\mathbf{K}_{e}=\sum_{C=1}^{n_{c}}A_{C}\,\widetilde{\mathbf{B}}^{(C)T}\mathbf{k}\,\widetilde{\mathbf{B}}^{(C)}.$ (19)
The capacity matrix is computed using the standard interpolation of $h$ as
$\mathbf{M}_{e}=\int_{\Omega_{e}}N^{T}S_{s}N\,d\Omega=\sum_{C=1}^{n_{c}}\int_{\Omega_{C}}N^{T}S_{s}N\,d\Omega,$ (20)
where the subcell integral may be evaluated by standard area quadrature on $\Omega_{C}$.
@@ 2.3 Element matrices with smoothed gradients
After assembly over all elements and enforcement of essential boundary conditions on $\Gamma_{h}$, the global system takes the form of Eq. (9) with $\mathbf{F}=\mathbf{f}_{p}+\mathbf{f}_{q}$.
2.4 Time integration
Let the time step be $\Delta t$ and denote the solution at $t^{n}$ by $\mathbf{H}^{n}$. Using the generalized $\theta$-method [37], the time derivative is approximated by
$\dot{\mathbf{H}}^{\,n+\theta}\approx\frac{\mathbf{H}^{n+1}-\mathbf{H}^{n}}{\Delta t},\qquad\mathbf{H}^{n+\theta}=(1-\theta)\mathbf{H}^{n}+\theta\mathbf{H}^{n+1},$ (23)
with $\theta\in(0,1]$. Evaluating $\mathbf{K}$, $\mathbf{M}$, and $\mathbf{F}$ at $t^{n+\theta}$ and applying the $\theta$-method to Eq. (9) yields
$\left(\frac{1}{\Delta t}\mathbf{M}^{n+\theta}+\theta\,\mathbf{K}^{n+\theta}\right)\mathbf{H}^{n+1}=\mathbf{F}^{n+\theta}+\left(\frac{1}{\Delta t}\mathbf{M}^{n+\theta}-(1-\theta)\,\mathbf{K}^{n+\theta}\right)\mathbf{H}^{n}.$ (24)
In particular, for the fully implicit backward Euler scheme ($\theta=1$), Eq. (24) reduces to
@@ 3 Polygonal interpolation using Wachspress basis functions
Consider a convex $n$-gon $\Omega_{e}$ with ordered vertices $\{\mathbf{x}_{k}\}_{k=1}^{n}$ in the physical plane. Fig. 1 provides a schematic illustration of the Wachspress shape functions on $\Omega_{e}$ [40] . Let $f_{1}$ and $f_{2}$ denote the two edges adjacent to vertex $k$ and let $\mathbf{n}_{f_{i}}$ be the corresponding outward unit normals. For a point $\mathbf{x}\in\Omega_{e}$, define the signed distance to the supporting line of edge $f_{i}$ as
$h_{f_{i}}(\mathbf{x})=\big(\mathbf{x}-\mathbf{x}_{f_{i}}\big)\cdot\mathbf{n}_{f_{i}},$ (26)
where $\mathbf{x}_{f_{i}}$ is any point on edge $f_{i}$ (e.g., one of its endpoints). With this convention, $h_{f_{i}}(\mathbf{x})>0$ for $\mathbf{x}$ inside $\Omega_{e}$.
Figure 1: Illustration of Wachspress shape functions on a convex polygon.
The Wachspress weight associated with vertex $k$ is then defined as
@@
where $\det(\mathbf{a},\mathbf{b})$ denotes the $2\times 2$ determinant. For a fixed polygon, $\det(\mathbf{n}_{f_{1}},\mathbf{n}_{f_{2}})$ is a constant that depends only on the local edge geometry at vertex $k$.
The Wachspress shape function associated with vertex $k$ is given by
@@
For convex polygons, the Wachspress basis is nonnegative, $C^{0}$-continuous, and linearly complete, and it satisfies the partition-of-unity property
@@
In the polygonal CSFEM, Wachspress functions enter the formulation only through boundary integrals of the form
@@
and their edge-wise representation
@@
which involve only $N_{I}$ but not $\nabla N_{I}$. Since $N_{I}$ is evaluated along straight edges, the integrand is smooth on each edge, and accurate results can be obtained using a small number of Gauss points.